OpenStudy (anonymous):

Evaluate the limit. As h-->0, lim (3+h)^2-3^2/h

OpenStudy (anonymous):

is \[\lim_{h \rightarrow 0}\left( 3+h \right)^{2}-3^{2/h}\] ????

OpenStudy (anonymous):

No, it's [(3+h)^2-3^2]/h...sorry.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}((3+h)^{2}-3^{2})/(h) = \lim_{h \rightarrow 0}(3+h-3)(3+h+3)/h\] \[\lim_{h \rightarrow 0}h(h+6)/h = \lim_{h \rightarrow 0}h+6=6\]

OpenStudy (anonymous):

remember \[a^{2}-b^{2} = (a+b)(a-b)\]